# Network-Optimised Spiking Neural Network for Event-Driven Networking

### v3

@@ Abstract
Time-critical networking requires low-latency decisions from sparse and bursty telemetry, where fixed-step neural inference waste computation. We introduce Network-Optimised Spiking (NOS)  [^1] https://mbilal84.github.io/nos-snn-networking/, a two-state neuron whose variables correspond to normalised queue occupancy and a recovery resource. NOS combines a saturating excitability nonlinearity for finite buffers, service and damping leaks, graph-local inputs with per-link gates and delays, and differentiable resets compatible with surrogate gradients and neuromorphic deployment. We establish existence and uniqueness of subthreshold equilibria, derive Jacobian-based local stability tests, and obtain a scalar network stability threshold that separates topology from node physics through a Perron-mode spectral condition. A stochastic arrival model aligned with telemetry smoothing links NOS responses to classical queueing behaviour while explaining increased variability near stability margins. Across chain, star, and scale-free graphs, NOS improves early-warning F1 and detection latency over MLP, RNN, GRU, and temporal-GNN baselines under a common residual-based protocol, while providing practical calibration and stability rules suited to resource-constrained networking deployments.
@@
Spiking neural networks, Congestion control, Active queue management, Explicit Congestion Notification, Network stability, Spectral graph theory, Distributed control, Queueing theory
@@ 1 Introduction
Machine learning has reshaped core networking tasks, including short-horizon traffic forecasting, anomaly detection, and routing optimisation. Conventional deep models such as multilayer perceptrons, recurrent networks, and graph neural networks perform well when labels are plentiful and training is done offline. Spiking neural networks (SNNs) offer a complementary path. Their event-driven computation is naturally sparse in time and energy, which aligns with packet-level operation at the edge and in devices with tight power budgets. Recent surveys and hardware results report large efficiency gains for spiking workloads on neuromorphic substrates, motivating designs that couple spiking dynamics to network semantics [Chowdhury2025, Davies2018, Tia2019]. Foundational work on temporal coding and computational power further motivates spike-based processing in control settings [IJCNN2000]. In parallel, hardware platforms such as Intel’s Loihi and its toolchain demonstrate on-chip learning and practical workflows for deploying SNNs [Davies2018, LoihiToolchain2018].
Classical SNNs, while powerful for neuroscientific exploration, are poorly matched to networking practice due to several fundamental limitations. Firstly, they often prioritize biological analogy, defining abstract internal states rather than variables with direct network interpretations like queue occupancy, service rate, or link delay [GUO2023] [Emre2019]. This biological focused modeling also leads to dynamics that present optimization challenges; the hard threshold-and-reset dynamics intrinsic to spiking neurons create a non-differentiable discontinuity that obstructs gradient-based optimisation, a cornerstone of modern machine learning [Abbott2016][Zenke2014]. Furthermore, the treatment of topology and per-link quality is often implicit or uniform, neglecting the complex, heterogeneous, and weighted graphs that define real-world networks. From a dynamical systems perspective, the excitability in common SNN models formulation can be unbounded, leading to unrealistic behaviour under high input load and presenting significant numerical instability [Yao2024]. Finally, implementations face practical limits in training cost, input encoding, and deployment on resource-constrained neuromorphic or edge hardware. Consequently, evaluation remains limited, often failing to bridge the gap between abstract simulations and performance on real network topologies and traffic patterns [GUO2023] [yang2022]. Classical spiking and queueing-inspired models, including Hodgkin–Huxley [Hausser2000], Izhikevich [Izhikevich2003] and G-networks [Gelenbe1994], either impose heavy dynamics or lack graph-aware, differentiable, queue-semantics. These gaps motivate a compact spiking unit whose state maps to normalised queue occupancy and recovery, whose inputs are graph-local with delays and optional per-link gates, and whose reset is differentiable for gradient-based optimisation [Zenke2014]. A second driver is deployment. Neuromorphic systems such as Intel’s Loihi families demonstrate on-chip learning rules, synaptic delays, and event-driven execution that can reduce energy for small edge workloads while keeping latency low [Davies2018]. Related SNN studies in robotics and generic complex-network settings illustrate broader interest in topology and control, but their objectives differ from packet dynamics and queue semantics considered here [ACM2025Path, ACMISL2024, AICI2025]. Such systems reinforce the case for spiking designs that expose measurable parameters and training handles, so that network operators can calibrate models from telemetry and run them under tight power envelopes. An extended comparison of when spiking neural networks are preferable to conventional ML and GNN models for networking tasks is given in methods A, while B reviews three classical spiking paradigms and evaluates their suitability for queue semantics, topology and trainability in packet networks.
@@
This article develops a Network-Optimised Spiking neuron (NOS) and examines its behaviour from single node to network scale. The model is compact and two-state, with each state and parameter tied to measurable network quantities. Queue occupancy serves as the fast state, and a slower recovery variable captures post-burst slowdown. Service appears as leak terms. Inputs are graph-local with options for per-link gating and fixed delay. A saturating nonlinearity prevents unbounded growth and represents finite buffers. Two differentiable reset strategies are introduced—an event-triggered exponential soft reset and a continuous pullback shaped by a sigmoid—both enabling gradient-based training while preserving spiking dynamics.
The analysis gives equilibrium existence and uniqueness, stability tests from the Jacobian, and a network threshold that scales with the Perron eigenvalue of the coupling matrix. Saturation enlarges the stable region by reducing the effective excitability slope. An operational margin links damping with offered load and provides a two-parameter contour for planning and capacity checks. Stochastic arrivals are represented as shot noise or compound Poisson input with calibrated scales. Practical guidance covers parameter initialisation from telemetry, surrogate-gradient training with a homotopy in reset sharpness, and deployment on neuromorphic platforms, are detailed in methods C, D and D.1, respectively. The evaluation protocol uses label-free forecasting with residual-based event scoring, in line with established practice in time-series anomaly detection.
Findings show that bounded excitability yields unique single-node equilibria and wider stability regions than the unbounded quadratic drive. Network stability follows the Perron eigenvalue, giving a direct means of control through weight normalisation and topology choice. Differentiable resets remove optimisation difficulties caused by hard jumps and enable surrogate-gradient training without disturbing event timing. In residual-based forecasting, NOS acts as a physics-aware forecaster, achieving competitive accuracy and latency on chain, star, and scale-free graphs while remaining interpretable and suitable for neuromorphic execution.
Results
@@ 2 Proposed model: Network-Optimised Spiking neural network (NOS)
where $S_{j}(t)$ is a presynaptic event train and $w_{ij}$ encodes link capacity, policy weight, or reliability. Three implementation choices motivate the NOS design. First, the reset must be continuous and differentiable so gradient-based training is possible. Second, state variables and parameters should map to observables such as normalised queue length, service rate, and recovery time. Third, subthreshold integration needs explicit stochastic drive so gradual load accumulation and burstiness are both represented. Implementation-level choices, detailed calibration rules, and additional parameter-tuning heuristics, including the pipeline in Fig. 6, are summarised in methods D.
@@ 2.1 Variable reinterpretation
[…]
@@ 2.2 State variables and dynamics
$\frac{dv_{i}}{dt}$ $=f_{\mathrm{sat}}(v_{i})+\beta v_{i}+\gamma-u_{i}+I_{i}(t)-\lambda v_{i}-\chi\bigl(v_{i}-v_{\mathrm{rest}}\bigr),$ (2)
$\frac{du_{i}}{dt}$ $=a\bigl(bv_{i}-u_{i}\bigr)-\mu u_{i}\;=\;ab\,v_{i}-(a+\mu)\,u_{i}.$ (3)
@@
Equation (2) separates four operational effects. The term $f_{\mathrm{sat}}(v_{i})$ (see (5)) captures the convex rise of backlog during rapid arrivals while remaining bounded. The linear part $\beta v_{i}+\gamma$ fits residual slope and offset that are visible in practice after removing coarse trends. The coupling $u_{i}$ models recovery drag that slows re-accumulation immediately after events. Finally, $\lambda v_{i}-\chi(v_{i}-v_{\mathrm{rest}})$ represents service and small-signal damping. Equation (3) integrates recent congestion with sensitivity $ab$ and relaxes at rate $a+\mu$. Thus, the NOS model reflects three intuitive aspects of queue behaviour: the rise under increasing load, the draining through service, and the short-lived slowdown that follows bursts. Fig. 1 illustrates a single NOS unit, highlighting its graph-local inputs, excitability dynamics, recovery variable, stochastic threshold, and differentiable reset pathway. Spike generation and reset principles are discussed in methods E. Operational interpretations and data-driven initialisation rules for all NOS parameters are summarised in Table 3 in methods C, with further design guidance in methods D.
@@ 3.2 Jacobian and linear stability
Linearising a NOS node about a subthreshold equilibrium $(v^{*},u^{*})$ produces a $2\times 2$ Jacobian whose entries depend on the small signal slope $f^{\prime}(v^{*})$ and the coefficients $(\lambda,\chi,a,b,\mu,\beta)$. Standard trace and determinant inequalities then yield local stability conditions that tie the damping provided by $\lambda$ and $\chi$ to the maximum admissible slope of $f_{\mathrm{sat}}$ (methods K.2).
@@ 3.3 Network coupling and global stability
A convenient scalar “operational margin” summarises these existence bounds in terms of subthreshold damping $\chi$ and the maximum steady input $I_{\max}=\max_{i}I_{i}^{*}$. Positive margin corresponds to safe operation with headroom for burstiness; margin near zero signals proximity to loss of equilibrium. The trade-off between damping and sustained load is illustrated in Fig. 2, while the full inequalities and derivations are given in methods K.3, and the associated bifurcation structure and parameter trends (saddle–node versus Hopf onsets) are summarised in methods L.
Figure 2: Operational margin $\Delta_{\mathrm{op}}(\chi,I_{\max})$ as a function of subthreshold damping $\chi$ and maximum steady input $I_{\max}=\max_{i}I_{i}^{*}$. The red contour marks $\Delta_{\mathrm{op}}=0$; points above this curve (larger $I_{\max}$ for the same $\chi$) lie beyond the small-signal existence bound. Parameters $\{\alpha,\kappa,\lambda,\beta,a,b,\mu,\gamma,v_{\mathrm{rest}}\}$ are held fixed as in the text. methods K.3 gives the analytical definition of $\Delta_{\mathrm{op}}$ and one-dimensional sweeps versus $I_{\max}$, together with sensitivity to $\rho(W)$ and $g$.
@@ 3.4 Block Jacobian, network spectrum, and stability threshold
Linearising the coupled NOS network about a subthreshold equilibrium yields a $2N\times 2N$ block Jacobian whose dominant network contribution is proportional to $g\,\rho(W)\,f^{\prime}_{\mathrm{sat}}(v^{*})$, where $g$ is a global coupling gain and $\rho(W)$ is the spectral radius of the coupling matrix. The analysis shows that the first loss of small-signal stability is well predicted by the scalar condition
$g\,\rho(W)\,f^{\prime}_{\mathrm{sat}}(v^{*})<\Lambda,$ (7)
with $\Lambda$ the net drain defined from the local coefficients (cf. (90)). Parameters are non-dimensionalised so that the stability proxy $g\,\rho(W)\,f^{\prime}_{\mathrm{sat}}(v^{*})<\Lambda$ and the operational margin $\Delta_{\mathrm{net}}$ are comparable across sites, as detailed in the methods N. This separates node physics from topology and matches the operational proxy used in Section 4. The full block Jacobian, Perron-mode reduction, numerical spectral validation, and Gershgorin-style bounds are given in methods K.4 and K.5, while methods L characterises the saddle–node and Hopf-onsets and shows that the instability threshold obeys the spectral collapse relation in (85).
@@ 4.2 Predictive baseline: comparison with machine learning models
These differences have direct operational meaning. High precision with weak recall misses incidents that matter. High recall with weak precision floods operators with alarms. Lower forecast error stabilises residual scales and avoids threshold drift. Lower latency preserves time for pacing, admission, and rerouting. Taken together, the figures show that NOS gives an effective operating point across scale-free, star, and chain graphs. It preserves recall without inflating false positives and it warns early enough to act when load begins to rise.
@@ Discussion
NOS treats congestion monitoring and control as a compact spiking process whose ingredients are aligned with how networks are actually run. The queue proxy $v$ is normalised to buffer fullness, the recovery state $u$ encodes recent stress, excitability is bounded to keep queues within a finite envelope, and the linear drains $\lambda$ and $\chi$ sit on the same device clocks as schedulers and EWMA filters. Thresholds implement budgeted alarms and resets are smooth enough for gradient-based training while still producing crisp, scheduler-scale events. Inputs are graph-local and carry explicit delays and per-link gates, so timing and residual capacity appear as first-class quantities rather than side effects of architecture.
@@
The comparison with classical spiking paradigms clarifies what is gained. Random Neural Networks exploit product forms and queueing intuition but hinge on Poisson assumptions and stationary averages. Hodgkin–Huxley and Izhikevich offer rich excitable dynamics, yet their variables and parameters are tied to ion channels and generic cortical behaviour rather than to queue occupancy, service, and link delay. Across these models, either the physics is faithful and heavy or it is light and disconnected from networking semantics. NOS occupies a middle ground. It keeps a two-dimensional state with a saturating excitability $f_{\mathrm{sat}}$ and a recovery loop, and it assigns queue-level meaning to every parameter that appears in the small-signal slope, stability tests, and reset dynamics. This mapping is the basis for the operational rules in the Online Methods.
@@
The analytical results show that stability can be expressed in terms that are familiar to operators. At node level, the Routh–Hurwitz conditions reduce to a balance between the steepest local slope of $f_{\mathrm{sat}}(v)$ and the effective drain $\Lambda=\lambda+\chi+ab/(a+\mu)-\beta$. At network level, coupling enters only through $g\,\rho(W)$, so the proxy $g\,\rho(W)\,f^{\prime}_{\mathrm{sat}}(v^{*})<\Lambda$ cleanly separates topology from node physics. In practice this means that one can reason about headroom using three kinds of lever: adjust the graph through $\rho(W)$ by reweighting or sparsifying heavy fan-in, adjust the gain $g$ on stressed regions through scheduling and policy, or adjust the local margin $\Lambda$ through service, damping, and recovery parameters. The bifurcation and synchrony analyses confirm that this scalar picture captures the onset of both saddle–node tipping and Hopf-like oscillations, and the finite-size experiments show that the resulting thresholds sharpen with network size.
@@
The queueing baselines anchor NOS against classical theory and device-level intuition. In open loop, light-load calibration aligns the NOS mean with the M/M/1 prediction, while the bounded excitability and soft pull-back cause earlier bending near saturation and much shorter tails under bursty MMPP arrivals. From a networking point of view, NOS trades a small increase in approximation error at high load for a systematic reduction in deep-queue probability, which is precisely where SLO violations and drop storms arise. In closed loop, NOS-based controllers produce marking signals that are decisive and low in jitter, avoiding both the noisy oscillations of raw queue marking and the lag introduced by heavy low-pass filtering. These behaviours are consistent with the analytical headroom conditions and with the linear decay times extracted from the local Jacobian.
@@
The forecasting experiments indicate that NOS is competitive with compact learned baselines when labels are scarce. A fluid predictor driven by the same arrivals gives a strong MAE reference, since it essentially reproduces the queueing simulator. NOS does not attempt to mimic every backlog micro-fluctuation and therefore carries a larger point error, yet it delivers high event skill for burst onset detection and produces compact responses around excursions. Temporal GNNs, MLPs and GRUs trained in the same label-free regime tend to trade off noise suppression against timing accuracy. In contrast, the bounded, reset-driven dynamics of NOS grant sharp onset timing without extended false positives between bursts. The protocol is deliberately strict, with train-only calibration and validation-based thresholds, so these gains are unlikely to be artefacts of tuning.
The noise sensitivity and avalanche analyses view NOS through the lens of stochastic drive. Shot-noise arrivals with realistic smoothing produce firing statistics whose dependence on amplitude, coupling and recovery matches the linear response theory. As the network approaches the coupling threshold, the same noise leads to higher firing rate, larger inter-spike-interval variability and heavier avalanche tails. The fitted power-law exponents and the synchrony order parameter show that coupling can push the system toward scale-free congestion cascades along dominant influence paths. From an operational perspective, this confirms that the same levers that increase deterministic headroom also reduce susceptibility to noise-driven cascades: higher service and damping, earlier saturation through larger $\kappa$, and reduced effective coupling on noisy links.
@@
A further aspect of the design is the explicit path to neuromorphic deployment. Address–event routing realises sparse, delayed coupling with the same $W$ and $\tau_{ij}$ that appear in the analysis. Soft resets are implemented as leaky pull-backs without algebraic jumps, and fixed-point quantisation respects the queue normalisation used in software. The stability proxy carries over once finite precision errors are bounded, so the same margin checks can be applied when deploying on mixed-signal or digital neuromorphic substrates. This alignment between model, analysis and hardware is attractive for edge and in-network deployments where energy per decision must scale with activity rather than with link count.

### v4

@@ Abstract
Delay-coupled systems often require low-latency decisions from sparse telemetry, where dense fixed-step neural inference is wasteful and can degrade near stability margins. We introduce Network-Optimised Spiking (NOS)  [^1] https://mbilal84.github.io/nos-snn-networking/, a trainable two-state event-driven dynamical unit for delayed, graph-coupled streams, whose states map to a fast load variable and a slower recovery resource. NOS uses bounded excitability for finite buffers, explicit leak terms for service and damping, and graph-local coupling with per-link gates and communication delays, with differentiable resets compatible with surrogate-gradient training and neuromorphic execution. We prove existence and uniqueness of subthreshold equilibria, derive Jacobian-based stability conditions, and obtain a scalar network stability threshold that separates topology from node dynamics via a Perron-mode spectral condition. A stochastic arrival model aligned with telemetry smoothing explains increased variability as systems approach stability boundaries. On delayed graph forecasting and early-warning tasks from queue telemetry, NOS improves detection F1 and detection latency over MLP, RNN/GRU, and temporal GNN baselines under a common residual-based protocol, while providing calibration rules for resource-constrained deployments.
@@
Event-driven computation, spiking neural networks, delay-coupled dynamical systems, resource-constrained inference, spectral stability, networked control, queueing dynamics, distributed systems
@@ 1 Introduction
Machine learning now underpins core networking tasks such as short-horizon traffic forecasting, anomaly detection, and routing optimisation. Standard deep models including multilayer perceptrons, recurrent networks, and graph neural networks are effective when training data are abundant and the inference budget can be treated as fixed. In operational networks, however, telemetry is often sparse and bursty, feedback is delayed, and performance can change sharply near stability margins. In this regime, dense fixed-step inference can waste computation and can also amplify timing errors, which matters when the target is early warning and safe operation rather than only average accuracy.
@@
Spiking neural networks (SNNs) offer a complementary computational regime. Their event-driven updates can be sparse in time and energy, which aligns with packet-level dynamics at the edge and with tight power envelopes. Surveys and neuromorphic results report substantial efficiency gains for spiking workloads on dedicated substrates, encouraging designs that connect spiking dynamics to system semantics rather than to abstract latent states [6, 7, 5]. Foundational work on temporal coding and the computational power of spikes also supports spike-based processing for time-critical inference and control [20]. In parallel, platforms such as Intel’s Loihi and its toolchain demonstrate practical workflows for SNN deployment, including on-chip learning and synaptic delays [7, 22].
Despite this promise, classical SNN formulations are often a poor fit for networking. Many models prioritise biological analogy and introduce internal states that do not correspond to measurable quantities such as queue occupancy, service rate, or link delay [14, 27]. Their hard threshold-and-reset mechanisms create non-differentiable discontinuities that complicate gradient-based optimisation, which remains the main workhorse of modern machine learning [1, 38]. Topology and per-link heterogeneity are also frequently handled implicitly or uniformly, even though real networks are weighted, directed, and delay-bearing. Finally, common excitable dynamics can be effectively unbounded under sustained drive, producing unrealistic behaviour and numerical instability under high load [37]. These issues are compounded by practical constraints on encoding, training cost, and deployment on resource-constrained neuromorphic or edge hardware, which helps explain why many evaluations remain disconnected from real topologies and traffic patterns [14, 36]. Related dynamical models including Hodgkin–Huxley [15], Izhikevich [17], and G-networks [10] either impose heavy dynamics or do not provide a graph-aware, differentiable design with explicit queue semantics.
We address these gaps by introducing Network-Optimised Spiking (NOS), a compact two-state event-driven unit designed for delayed, graph-coupled telemetry streams in networks. NOS maps its state directly to normalised queue occupancy and a recovery resource, uses bounded excitability to respect finite buffers, includes explicit leak terms for service and damping, and supports graph-local coupling with optional per-link gates and communication delays. Crucially, NOS replaces hard resets with differentiable reset mechanisms that preserve event timing while enabling surrogate-gradient training [38]. We also motivate deployment on neuromorphic substrates where synaptic delays and event-driven execution are native [7]. Related SNN studies in robotics and complex-network settings consider topology and control, but their objectives differ from packet dynamics and queue semantics, and they do not provide the calibration hooks needed for network operation [39, 32, 21]. An extended discussion of when SNNs are preferable to conventional deep and graph models for networking tasks is given in Section A, while Section B reviews classical spiking paradigms and evaluates their suitability for queue semantics, topology, and trainability.
@@
(i) A semantics-aligned spiking unit: we propose a two-state NOS model whose variables and parameters map to measurable network quantities, with bounded excitability, explicit service and damping leaks, and graph-local delayed coupling with optional per-link gating. (ii) Trainability via differentiable resets: we introduce two differentiable reset strategies, an event-triggered exponential soft reset and a continuous pullback shaped by a sigmoid, to support surrogate-gradient optimisation while preserving spiking behaviour. (iii) Stability analysis at node and network scale: we prove existence and uniqueness of subthreshold equilibria, derive Jacobian-based local stability tests, and obtain a scalar network stability threshold that scales with the Perron eigenvalue of the coupling matrix; saturation enlarges the stable region by reducing the effective excitability slope. (iv) Operational guidance: we derive an operational margin that links damping to offered load, yielding a simple two-parameter contour for planning and capacity checks, and we provide calibration rules intended for delay-sensitive, resource-constrained deployments.
@@ 2 Proposed model: Network-Optimised Spiking neural network (NOS)
where $S_{j}(t)$ is a presynaptic event train and $w_{ij}$ encodes link capacity, policy weight, or reliability. Three implementation choices motivate the NOS design. First, the reset must be continuous and differentiable so gradient-based training is possible. Second, state variables and parameters should map to observables such as normalised queue length, service rate, and recovery time. Third, subthreshold integration needs explicit stochastic drive so gradual load accumulation and burstiness are both represented.
Implementation-level choices, detailed calibration rules, and additional parameter-tuning heuristics, including the pipeline in Fig. 6, are summarised in methods D.
@@ 2.1 Variable reinterpretation
vnormalised queue length or congestion $elv\;\mapsto\;\text${normalised queue length or congestion level}
$u\;\mapsto\;\text{recovery or slowdown resource}.$
@@ 2.2 State variables and dynamics
$\begin{split}\frac{dv_{i}}{dt}=f_{\mathrm{sat}}(v_{i})+\beta v_{i}+\gamma-\xi\,u_{i}+I_{i}(t)-\lambda v_{i}-\\ \chi\bigl(v_{i}-v_{\mathrm{rest}}\bigr),\end{split}$ (2)
$\begin{split}\frac{du_{i}}{dt}&=a\bigl(bv_{i}-u_{i}\bigr)-\mu u_{i}\;=\;ab\,v_{i}-(a+\mu)\,u_{i}.\end{split}$ (3)
Here $v$ is the normalised queue level (dimensionless). The recovery state $u$ is also taken dimensionless, and the scalar $\xi>0$ maps recovery into a rate contribution in (2) (so $\xi$ has units of inverse time, or per-bin under a discrete time base). In all experiments we absorb this scaling into the definition of $u$ and fix $\xi=1$, so $\xi$ is omitted from subsequent notation. Throughout, time $t$ is measured in sampling bins of width $\Delta t$ (default $\Delta t=5\,\mathrm{ms}$). Derivatives $d(\cdot)/dt$ are therefore per-bin rates, and all coefficients in (2)–(3) are dimensionless per-bin quantities. Physical rates in $\mathrm{s}^{-1}$ are obtained by $r_{\mathrm{phys}}=\frac{r}{\Delta t}.$
@@
Equation (2) separates four operational effects. The term $f_{\mathrm{sat}}(v_{i})$ (see (5)) captures the convex rise of backlog during rapid arrivals while remaining bounded. The linear part $\beta v_{i}+\gamma$ fits residual slope and offset that are visible in practice after removing coarse trends. The coupling $\xi u_{i}$ models recovery drag that slows re-accumulation immediately after events. Finally, $\lambda v_{i}-\chi(v_{i}-v_{\mathrm{rest}})$ represents service and small-signal damping. Equation (3) integrates recent congestion with sensitivity $ab$ and relaxes at rate $a+\mu$. Thus, the NOS model reflects three intuitive aspects of queue behaviour: the rise under increasing load, the draining through service, and the short-lived slowdown that follows bursts. Fig. 1 illustrates a single NOS unit, highlighting its graph-local inputs, excitability dynamics, recovery variable, stochastic threshold, and differentiable reset pathway. Spike generation and reset principles are discussed in methods E. Operational interpretations and data-driven initialisation rules for all NOS parameters are summarised in Table 3 in methods C, with further design guidance in methods D.
@@ 3.2 Jacobian and linear stability
Linearising a NOS node about a subthreshold equilibrium $(v^{*},u^{*})$ produces a $2\times 2$ Jacobian whose entries depend on the small signal slope $f^{\prime}(v^{*})$ and the coefficients $(\lambda,\chi,a,b,\mu,\beta)$. Standard trace and determinant inequalities then yield local stability conditions that tie the damping provided by $\lambda$ and $\chi$ to the maximum admissible slope of $f_{\mathrm{sat}}$ (methods K.3).
@@ 3.3 Network coupling and global stability
A convenient scalar “operational margin” summarises these existence bounds in terms of subthreshold damping $\chi$ and the maximum steady input $I_{\max}=\max_{i}I_{i}^{*}$. Positive margin corresponds to safe operation with headroom for burstiness; margin near zero signals proximity to loss of equilibrium. The trade-off between damping and sustained load is illustrated in Fig. 2, while the full inequalities and derivations are given in methods K.5, and the associated bifurcation structure and parameter trends (saddle–node versus Hopf onsets) are summarised in methods L.
Figure 2: Operational margin $\Delta_{\mathrm{op}}(\chi,I_{\max})$ as a function of subthreshold damping $\chi$ and maximum steady input $I_{\max}=\max_{i}I_{i}^{*}$. The red contour marks $\Delta_{\mathrm{op}}=0$; points above this curve (larger $I_{\max}$ for the same $\chi$) lie beyond the small-signal existence bound. Parameters $\{\alpha,\kappa,\lambda,\beta,a,b,\mu,\gamma,v_{\mathrm{rest}}\}$ are held fixed as in the text.
@@ 3.4 Block Jacobian, network spectrum, and stability threshold
Linearising the coupled NOS network about a subthreshold equilibrium yields a $2N\times 2N$ block Jacobian, whose network contribution enters through the linearised drive sensitivity $\delta I\approx G\,\delta v$. Under homogeneous scaling, one may take $G=gW$. A corresponding vector-norm bound is given in methods K.2. Projecting onto the Perron mode of $W$ reduces the stability check to a $2\times 2$ effective Jacobian (methods K.6). Routh–Hurwitz then yields a scalar coupling threshold in terms of $k:=g\,\rho(W)$:
$g\,\rho(W)\;<\;\Lambda-f^{\prime}_{\mathrm{sat}}(v^{*}),$ (7)
equivalently $g\,\rho(W)+f^{\prime}_{\mathrm{sat}}(v^{*})<\Lambda$, where $\Lambda$ is the net drain defined from the local coefficients (cf. (105)). Parameters are non-dimensionalised so that $T\,\Lambda$ and $T\,f^{\prime}_{\mathrm{sat}}(v^{*})$ are dimensionless and the operational margin $\Delta_{\mathrm{net}}:=\Lambda-f^{\prime}_{\mathrm{sat}}(v^{*})-g\,\rho(W).$ is comparable across sites (methods N). This separates node physics from topology and matches the operational proxy used in Section 4.
The full block Jacobian, Perron-mode reduction, numerical spectral validation, and Gershgorin-style bounds are given in methods K.6 and K.7. Finally, methods L characterises the saddle–node and Hopf onsets and shows that the instability threshold obeys the spectral collapse relation in (100).
@@ 5 Discussion
NOS treats congestion monitoring and control as a compact spiking process whose ingredients reflect how networks are actually run. The queue proxy $v$ is normalised to buffer fullness, the recovery state $u$ encodes recent stress, excitability is bounded to keep queues within a finite envelope, and the linear drains $\lambda$ and $\chi$ sit on the same device clocks as schedulers and EWMA filters. Thresholds implement budgeted alarms and resets are smooth enough for gradient-based training while still producing crisp, scheduler-scale events. Inputs are graph-local and include explicit delays and per-link gates, so timing and residual capacity appear as first-class quantities rather than side effects of architecture.
@@
The comparison with classical spiking paradigms clarifies what is gained. Random Neural Networks exploit product forms and queueing intuition but hinge on Poisson assumptions and stationary averages. Hodgkin–Huxley and Izhikevich offer rich excitable dynamics, yet their variables and parameters are tied to ion channels and generic cortical behaviour rather than to queue occupancy, service, and link delay. Across these models, either the physics is faithful and heavy or it is light and disconnected from networking semantics. NOS occupies a middle ground. It keeps a two-dimensional state with saturating excitability $f_{\mathrm{sat}}$ and a recovery loop, and it assigns queue-level meaning to every parameter that appears in the small-signal slope, stability tests, and reset dynamics. This mapping underpins the operational rules in the methods.
@@
The analytical results show that stability can be expressed in terms that are familiar to operators. At node level, the Routh–Hurwitz conditions reduce to a balance between the local slope $f^{\prime}_{\mathrm{sat}}(v^{*})$ and a drain margin $\Lambda:=\min\{(a+\mu),\,ab/(a+\mu)\}-(\beta-\lambda-\chi)$. At network level, coupling enters through $k=g\rho(W)$, giving the proxy $k+f^{\prime}_{\mathrm{sat}}(v^{*})<\Lambda$. At network level, coupling enters only through $g\,\rho(W)$, so the proxy $g\,\rho(W)\,+f^{\prime}_{\mathrm{sat}}(v^{*})<\Lambda$ separates topology from node physics. This yields three practical levers: adjust $\rho(W)$ by reweighting or sparsifying heavy fan-in, adjust the gain $g$ on stressed regions through scheduling and policy, or adjust the local margin $\Lambda$ through service, damping, and recovery parameters. The bifurcation and synchrony analyses support this scalar picture, capturing both saddle–node tipping and Hopf-like oscillations, while finite-size experiments show that thresholds sharpen with network size.
@@
Queueing baselines anchor NOS against classical theory and device-level intuition. In open loop, light-load calibration aligns the NOS mean with the M/M/1 prediction, while bounded excitability and soft pull-back bend earlier near saturation and yield shorter tails under bursty MMPP arrivals. From a networking viewpoint, NOS trades a small increase in point error at high load for a systematic reduction in deep-queue probability, which is where SLO violations and drop storms arise. In closed loop, NOS-based controllers produce marking signals that are decisive and low in jitter, avoiding both noisy oscillations of raw queue marking and the lag induced by heavy low-pass filtering. These behaviours are consistent with the headroom conditions and with the linear decay times extracted from the local Jacobian.
@@
The forecasting experiments indicate that NOS is competitive with compact learned baselines when labels are scarce. A fluid predictor driven by the same arrivals provides a strong MAE reference because it effectively reproduces the queueing simulator. NOS does not attempt to mimic every backlog micro-fluctuation and therefore carries a larger point error, yet it delivers high event skill for burst onset detection and produces compact responses around excursions. Temporal GNNs, MLPs, and GRUs trained in the same label-free regime tend to trade off noise suppression against timing accuracy. In contrast, the bounded, reset-driven dynamics of NOS yield sharp onset timing without prolonged false positives between bursts. The protocol is deliberately strict, with train-only calibration and validation-based thresholds, which reduces the likelihood that gains arise from tuning artefacts. Noise sensitivity and avalanche analyses view NOS through the lens of stochastic drive. Shot-noise arrivals with realistic smoothing produce firing statistics whose dependence on amplitude, coupling, and recovery matches the linear response theory. As the network approaches the coupling threshold, the same noise induces higher firing rates, larger inter-spike-interval variability, and heavier avalanche tails. The fitted exponents and synchrony measures indicate that coupling can push the system toward congestion cascades along dominant influence paths. Operationally, this supports the same levers suggested by the deterministic analysis: increase service and damping, saturate earlier through larger $\kappa$, and reduce effective coupling on noisy links.
@@
An explicit path to neuromorphic deployment follows from the design. Address–event routing realises sparse, delayed coupling with the same $W$ and $\tau_{ij}$ that appear in the analysis. Soft resets can be implemented as leaky pull-backs without algebraic jumps, and fixed-point quantisation respects the queue normalisation used in software. The stability proxy carries over once finite-precision errors are bounded, allowing the same margin checks to be applied on mixed-signal or digital neuromorphic substrates. This alignment between model, analysis, and hardware is attractive for edge and in-network deployments where energy per decision should scale with activity rather than with link count.